# ToolRM: Outcome Reward Models for Tool-Calling Large Language Models

### v1

@@ 5 Results
RQ2: Can ToolRM improve the performance during inference through Best-of-$n$ sampling?, and
RQ3: When ToolRM is used for filtering training data, can it lead to improved finetuned models?
@@ 5.1 RQ1: FC-RewardBench evaluation
We evaluate ToolRM against state-of-the-art RMs from RewardBench (Lambert et al., 2024), as well as leading LLMs used in an LLM-as-a-Judge setting, on the FC-RewardBench dataset.
RMs are evaluated by comparing scores assigned to the correct tool call outputs and incorrect tool call outputs for the same input. A prediction is counted as correct when the score for the correct tool call exceeds that of the incorrect one. LLMs-as-Judges are evaluated with a pairwise comparison prompt, where both candidate tool calls are presented and the model is instructed to select the correct one. To avoid position bias, the order of candidates is randomized. Experimental details, including the full prompt template, are provided in Appendix A.4.
We show the results in Figure 1 and observe the following:
  • Specialized RMs under-perform on tool-calling tasks. Despite strong results on RewardBench domains, most specialized RMs fail to generalize effectively to the tool-calling domain. While some individual variants achieve higher scores, performance remains inconsistent and generally below state-of-the-art performance.
  • LLMs-as-Judges achieve higher accuracy but are computationally expensive. LLMs-as-Judges attain strong performance on FC-RewardBench (exceeding 80% across all models), but their large parameter counts impose substantial computational costs.
  • ToolRM achieves the highest accuracy on the benchmark while maintaining efficiency with respect to model size. The ToolRM-14B and ToolRM-7B variants outperform all other generative and sequential classifier models. Notably, even the ToolRM-1.5B variant surpasses the gpt-oss-120B model, approaching the performance of substantially larger Llama-4 models.
Figure 3: Correlation heatmap between performance on FC-RewardBench and downstream accuracy across generator models and benchmarks, showing consistently strong alignment (avg. correlation = 0.84).
Correlation with performance on downstream tasks:
Figure 4: Performance of the Qwen3 series (top) and xLAM-2 series (bottom) in the Best-of-$n$ $(n=32)$ setting across five function-calling benchmarks: API-Bank-1, API-Bank-2, NexusRaven, ToolAlpaca, and SealTools.
@@ 5.2 RQ2: Best-of-$n$ sampling with ToolRM
Qwen3-14b Greedy 68.62 89.25 82.16 33.75
ToolRM-7B 68.29 91.88 84.16 32.50
ToolRM-14B 69.18 92.08 83.57 34.25
Qwen3-32b Greedy 69.19 89.33 82.83 38.25
ToolRM-7B 68.33 90.75 84.90 35.25
ToolRM-14B 70.61 92.31 84.23 39.62
@@
BFCL V3 ToolAlpaca Nexus API-Bank-1 API-Bank-2 Sealtools AVG
Llama-3.1-8B-Instruct 49.6 38.0 64.8 67.9 66.2 37.6 54.0
Llama-3.1-8B-Instruct (tuned - 16K) 54.1 43.0 75.5 57.4 63.5 72.7 61.0
Llama-3.1-8B-Instruct (tuned - random 8K) 55.2 44.0 74.2 49.9 54.1 73.2 58.4
Llama-3.1-8B-Instruct (tuned - best scored 8K) 55.4 44.0 72.0 63.7 66.2 73.7 62.5
Table 3: Finetuning results of Llama-3.1-8B-Instruct on three training subsets: full 16K dataset (tuned – 16K), 8K randomly sampled subset (tuned – random 8K), and top 8K examples selected via ToolRM (tuned – best scored 8K).
In this section, we evaluate ToolRM in a Best-of-$n$ setting across multiple generator models. For each input, we sample $n=32$ independent generations from the generator model and use ToolRM to score and select the highest-ranked generation as the final output. Intuitively, a stronger RM should more reliably identify the correct tool call, thereby improving task performance. We compare against two baselines: Greedy Decoding and Majority Voting ($n=32$), where the most frequently occurring final answer is selected as the final output. For non-BFCL benchmarks, we report the Full Sequence Matching metric Basu et al. (2025), which checks whether the predicted tool sequence—including tool names and argument-value pairs—exactly matches the gold sequence. For BFCL, we use its native evaluation metrics: AST-based scores for single-turn tasks and state-based/response-based metrics for multi-turn cases.
Figure 4 reports average performance across five benchmarks (API-Bank-Level-1, API-Bank-Level-2, ToolAlpaca, NexusRaven, and SealTools), while Table 2 presents results on the BFCL-v3 dataset. We summarize the key insights below:
@@
  • Small Language Models (SLMs) benefit the most: Best-of-$n$ sampling with Qwen3-0.6B and ToolRM-14B as the ranker improves accuracy from 39.5% to 64.38% – a gain of 24.9 points on non-BFCL benchmarks (Figure 4). This performance surpasses that of Qwen3-32B (63.8%) and Llama-xLAM-2-70b-fc-r (63.6%) with greedy decoding. On BFCL-v3, xLAM-2-1B-fc-r with ToolRM-14B improves overall, Non-Live AST, and Live AST accuracies by 3.2, 6.5, and 6.2 points, respectively. Qwen3-1.7B achieves even larger improvements of 5.3, 9.6, and 8.7 points on these metrics (Table 2).
  • SLM + RM can match or surpass larger models: Best-of-$n$ sampling with Qwen3-8B and ToolRM-14B improves non-BFCL benchmark accuracy by 6.8% points to 70.48%, which is 5.6 points higher than the best greedy baseline. On BFCL-v3, the same setup yields gains of 3.3 points on Non-Live AST and 2.9 points on Live AST, exceeding the performance of Qwen3-32B with greedy decoding.
  • RM scale correlates with gain: We present Best-of-$n$ results with ToolRM-1.5B, 7B, and 14B in Figure 4 and results with ToolRM-7B and 14B in Table 2. In nearly all cases, Best-of-$n$ performance correlates with ToolRM size; larger RMs lead to better performance improvements. We do note, however, that the smallest ToolRM-1.5B model provides substantial gains and can be useful in compute-constrained scenarios.
  • Gains are consistent across benchmarks and model families: Although SLMs exhibit the largest improvements, we observe consistent gains across (a) all considered benchmarks, (b) generator model scales ranging from 0.6B to 70B, and (c) both general instruction-tuned and function-calling–specialized models.
@@
  • Diminishing returns for very large models: Improvements for large-scale generators (32B+) are modest. For instance, Llama-xLAM-2-32B-fc-r improves by only 2.1 points on non-BFCL benchmarks and 2.5 points on BFCL Live AST accuracy, suggesting limited additional utility of Best-of-$n$ sampling with very strong base models.
@@
In Appendix A.5, we provide a breakdown of errors with greedy decoding and with Best-of-$n$ sampling with ToolRM-14B.
5.3 RQ3: ToolRM For Data Filtering
@@
In this experiment, we assess the effectiveness of using ToolRM as a data filter to construct a high-quality training dataset for tool-use models. We curate a training corpus comprising both single-turn and multi-turn examples drawn from APIGen-MT Liu et al. (2024c), SealTools Wu et al. (2024), Glaive V2999https://huggingface.co/datasets/glaiveai/glaive-function-calling-v2, and Granite function-calling dataset Abdelaziz et al. (2024), yielding a total of 16K samples. We highlight that the data sources selected for this experiment have no overlap with ToolRM training data, thus allowing us to test the generalization capabilities of ToolRM. We select Llama-3.1-8B-Instruct Grattafiori et al. (2024) as the base model and performed LoRA-based fine-tuning Hu et al. (2022) to train each variant for 1 epoch with a learning rate of 2e-4, a LoRA rank of 16, alpha of 32, a cosine scheduler, and a warmup ratio of 10%.
@@
Table 3 compares the performance of the base model with three fine-tuned variants: (1) trained on the full 16K dataset, (2) trained on a random 8K subset, and (3) trained on the top 8K samples as ranked by ToolRM-14B. We highlight the following key insights from the results.
@@
  • Fine-tuning consistently improves performance. All fine-tuned models outperform the base model, raising average accuracy from 54.0% to 61.0% when trained on the full dataset.
  • Naive subsampling degrades performance. Training on a random 8K subset reduces accuracy to 58.4% – 2.6 points below the model fine-tuned on the entire corpus. This performance degradation reflects the inapplicability of naively subsampling to reduce dataset size since it results in both the inclusion of low-quality samples and the exclusion of high-quality ones.
@@
  • ToolRM based data filtering achieves the best performance. Selecting the top 50% of samples with ToolRM-14B yields the strongest results at 62.5%, surpassing the full-data model while using only half the training corpus. This demonstrates that ToolRM can effectively identify high-quality data, enabling superior performance under tighter data budgets.
@@
Overall, these results highlight the importance of data quality in fine-tuning tool-use models and show that filtering low-quality samples through a reward model can yield superior performance while reducing dataset size.
@@ 6 Conclusion
In this paper, we presented a comprehensive framework for reward modeling in tool-calling scenarios, addressing a critical gap in current LLM evaluation pipelines. Our benchmark, FC-RewardBench, enables systematic assessment of reward models on tool-based reasoning. We also presented a framework for training outcome RMs that outperform existing significantly larger RMs in the tool calling setting. When used for inference-time scaling, our RMs also improve the performance of general-purpose and tool calling models by up to 25% on various tool calling benchmarks. Looking ahead, we see several promising directions for advancing reward modeling in this domain. First, moving beyond classification-based RMs to generative verifiers with chain-of-thought reasoning could improve robustness and interpretability. Second, incorporating tool and environment state into training could help models safely recover from execution failures. Finally, bridging outcome and process reward modeling may offer a unified framework that balances scalability with fine-grained control over reasoning quality.
@@ A.4.2 LLM-as-Judge Prompt
A.5 Error Analysis

### v2

@@ 5 Results
RQ2: Can ToolRM improve the performance during inference through Best-of-$n$ sampling? And,
RQ3: Can ToolRM lead to better models through reward-guided data filtering or policy optimization?
@@ 5.1 RQ1: FC-RewardBench evaluation
We evaluate ToolRM against state-of-the-art RMs from RewardBench (Lambert et al., 2024), Tool-Augmented RM (Themis) (Li et al., 2024), as well as leading LLMs used in an LLM-as-a-Judge setting, on the FC-RewardBench dataset.
RMs are evaluated by comparing scores assigned to the correct tool call outputs and incorrect tool call outputs for the same input. A prediction is counted as correct when the score for the correct tool call exceeds that of the incorrect one. LLMs-as-Judges are evaluated with a pairwise comparison prompt, where both candidate tool calls are presented and the model is instructed to select the correct one. To avoid position bias, the order of candidates is randomized. Experimental details, including the full prompt template, are provided in Appendix A.4. We show the results in Figure 1 and observe the following:
  • Existing reward models struggle with tool-calling tasks. Specialized RMs underperform and often fail to generalize to tool-calling behaviors – e.g., the Tool-Augmented RM (Themis) achieves only 45% accuracy on FC-RewardBench, and rule-based method fares even worse. While LLMs-as-Judges achieve strong accuracy (exceeding 80%), their large parameter counts make them computationally expensive.
  • ToolRM achieves state-of-the-art accuracy with high efficiency. ToolRM-7B and ToolRM-14B outperform all other generative and sequential classifier models, and the ToolRM-1.5B variant even surpasses the gpt-oss-120B model, approaching the performance of much larger Llama-4 models.
The primary purpose of FC-RewardBench is to enable quick evaluation of RMs without requiring costly downstream experiments. Therefore, performance on FC-RewardBench should correlate strongly with downstream results. In Appendix A.5, we report correlations between 11 RMs on FC-RewardBench and their Best-of-nnscores on five downstream benchmarks. With an average correlation of 0.84, FC-RewardBench provides a reliable and efficient proxy for downstream RM evaluation.
@@ 5.2 RQ2: Best-of-$n$ sampling with ToolRM
In this section, we evaluate ToolRM in a Best-of-$n$ setting across multiple generator models. For each input, we sample $n=32$ independent generations using temperature $T=0.6$ from the generator model and use ToolRM to score and select the highest-ranked generation as the final output. Intuitively, a stronger RM should more reliably identify the correct tool call, thereby improving task performance. We compare against three baselines: 1) Greedy Decoding, 2) Majority Voting – where the most frequently occurring final answer is selected as the output, and 3) Schema Validation – where we compare the output against the input tool schema and return the generation with the highest likelihood that validates the schema. For non-BFCL benchmarks, we report the Full Sequence Matching metric (Basu et al., 2025), which checks whether the predicted tool sequence – including tool names and argument-value pairs – exactly matches the gold sequence. For BFCL, we use its native evaluation metrics: AST-based scores for single-turn tasks and state-based/response-based metrics for multi-turn cases.
Figure 2 reports average performance across five benchmarks (API-Bank-1, API-Bank-2, ToolAlpaca, NexusRaven, and SealTools), while Table 2 presents results on the BFCL-v3 dataset. We summarize the key insights below.
Figure 2: Performance of the Qwen3 series and xLAM-2 series in the Best-of-$n$ $(n=32)$ setting across five benchmarks: API-Bank-1, API-Bank-2, NexusRaven, ToolAlpaca, and SealTools.
@@
  • Small Language Models (SLMs) benefit most and can match or surpass larger models: Best-of-$n$ sampling with ToolRM-14B yields the largest gains for small generators. Qwen3-0.6B improves from 39.5% to 64.4% – a gain of 24.9 points that surpasses Qwen3-32B (63.8%) and xLAM-2-70B (63.6%) with greedy decoding. Similarly, Qwen3-8B reaches 70.5%, exceeding all greedy baselines by 5.6 points. On BFCL-v3, Qwen3-1.7B achieves improvements of 5.3 points on overall accuracy and 9.6 and 8.7 points on Non-Live AST and Live AST metrics respectively (Table 2).
@@
  • Diminishing returns for very large models: Improvements for 32B+ generators are modest – Llama-xLAM-2-32B-fc-r gains 2.1 points on non-BFCL benchmarks and 2.5 points on BFCL Live AST – suggesting limited additional utility of Best-of-$n$ sampling with already-strong base models.
@@
We also look at the breakdown of errors with greedy decoding and Best-of-$n$ sampling with ToolRM-14B, and present the results in Appendix A.6.
Best-of-$n$ sampling improves model robustness:
We examine the impact of Best-of-$n$ sampling on model robustness to noise in the input. We utilize RoTBench (Ye et al., 2024), which comprises of 568 tool specifications and 105 user queries paired with tools with varying levels of noise. The Clean split contains tool and parameter names that clearly reflect their usage, while the Slight, Medium, and Heavy splits introduce increasing noise through operations such as character insertion and deletion, name reversal, and name swapping. The Union split combines all noisy variants and represents the most challenging setting. Model performance is evaluated across three tasks: Tool Selection, Parameter Identification, and Content Filling.
Generator Model Tool Selection Parameter Identification Content Filling
Clean Clean@32 Union Union@32 Clean Clean@32 Union Union@32 Clean Clean@32 Union Union@32
Qwen-1.7B 54.3 76.2 47.6 58.1 37.1 55.2 27.6 40.0 27.6 41.0 21.0 30.5
Qwen-8B 52.4 72.4 45.7 66.7 38.1 55.2 30.5 46.7 27.6 42.9 20.0 31.4
Qwen-32B 65.7 76.2 52.4 72.4 38.1 56.2 30.5 44.8 25.7 42.9 21.0 31.4
Table 3: Performance of Qwen models on RoTBench with greedy decoding (Clean and Union) and Best-of-$n$ ($n=32$) with ToolRM-14B (Clean@32 and Union@32).
Table 3 reports results for greedy decoding and Best-of-$n$ ($n=32$) with ToolRM-14B. We highlight two key findings. First, Best-of-$n$ decoding yields substantial gains across all models and tasks. For instance, ToolRM improved Qwen-8B performance on Tool Selection from 52.4 to 72.4 on the Clean split, while performance on the Union split improved from 45.7 to 66.7. Comparable gains of 15–25 points are observed for Parameter Identification and Content Filling. Second, Union@32 consistently outperforms the Clean baseline, despite Union being the more difficult split. For example, Qwen-32B achieves 72.4 on Tool Selection under Union@32 compared to 65.7 under Clean, showing that Best-of-$n$ decoding not only mitigates noise but can also exceed performance on noise-free data.
5.3 RQ3: Reward-guided Fine-Tuning
5.3.1 ToolRM for data filtering
@@
In this experiment, we assess the effectiveness of using ToolRM as a data filter to construct a high-quality training dataset for tool-use models. We curate a training corpus comprising both single-turn and multi-turn examples drawn from APIGen-MT (Liu et al., 2024c), SealTools (Wu et al., 2024), Glaive V2444https://huggingface.co/datasets/glaiveai/glaive-function-calling-v2, and Granite function-calling dataset (Abdelaziz et al., 2024), yielding a total of 16K samples. We highlight that these datasets have no overlap with ToolRM training data, thus allowing us to test the generalization capabilities of ToolRM. We select Llama-3.1-8B-Instruct (Grattafiori et al., 2024) as the base model and performed LoRA-based fine-tuning (Hu et al., 2022) to train each variant for 1 epoch with a learning rate of 2e-4, a LoRA rank of 16, alpha of 32, a cosine scheduler, and a warmup ratio of 10%.
Llama-3.1-8B-Instruct BFCL V3 ToolAlpaca Nexus API-Bank-1 API-Bank-2 Sealtools Average
Base 49.6 38.0 64.8 67.9 66.2 37.6 54.0
FT-16K 54.1 43.0 75.5 57.4 63.5 72.7 61.0
FT-Random-8K 55.2 44.0 74.2 49.9 54.1 73.2 58.4
FT-Best-8K 55.4 44.0 72.0 63.7 66.2 73.7 62.5
Table 4: Finetuning results of Llama-3.1-8B-Instruct on three training subsets: full 16K dataset (FT-16K), 8K randomly sampled (FT-Random-8K), and top 8K selected by ToolRM-14B (FT-Best-8K).
@@
Table 4 compares the performance of the base model with three fine-tuned variants: (1) trained on the full 16K dataset (FT-16K), (2) trained on a random 8K subset (FT-Random-8K), and (3) trained on the top 8K samples as ranked by ToolRM-14B (FT-Best-8K). We highlight the following key insights from the results.
@@
  • Fine-tuning improves performance, but naive subsampling degrades it. All fine-tuned models outperform the base model, increasing accuracy from 54.0% to 61.0% when trained on the full dataset. However, training on a random 8K subset drops accuracy to 58.4%, showing that naive subsampling includes low-quality samples and discards high-quality ones.
@@
  • ToolRM-based data filtering achieves the best results. Selecting the top 50% of samples using ToolRM-14B yields 62.5% accuracy – surpassing the full-data model while using only half the corpus – demonstrating ToolRM’s ability to identify high-quality data, enabling superior performance under tighter training budgets.
@@
These results highlight the importance of data quality in fine-tuning tool-use models and show that reward-guided filtering of low-quality data can yield superior performance with less training.
5.3.2 Policy Optimization Using ToolRM
To assess the utility of ToolRM for policy optimization, we follow ToolRL (Qian et al., 2025a) and train models using Group Relative Policy Optimization (GRPO) (Shao et al., 2024). ToolRL defines the reward as $R=R_{\text{format}}+R_{\text{correctness}}$, where $R_{\text{format}}\in\{0,1\}$ evaluates whether the output adheres to the format specified in the prompt, and $R_{\text{correctness}}$ measures the correctness of the tool-call output. We consider three variants of $R_{\text{correctness}}$: (1) $R_{\text{schema}}\in\{-1,1\}$, validates the predicted tool calls against the provided tool specifications, assigning 1-1 if there are schema violations and +1+1 otherwise; (2) $R_{\text{ToolRL}}\in[-3,3]$, follows Qian et al. (2025a) and computes rewards by comparing predicted and ground-truth tool calls; and (3) $R_{\text{ToolRM}}\in[-3,3]$, scores the tool calls using ToolRM-14B. Notably, $R_{\text{schema}}$ and $R_{\text{ToolRM}}$ do not require access to ground-truth tool calls, making them more appropriate for RL settings, whereas $R_{\text{ToolRL}}$ requires ground truth, limiting its applicability.
Model Reward Variant Non-Live AST Acc Live AST Acc
Llama-3.2- 3B-Instruct Base 15.35% 43.82%
$R_{\text{Schema}}$ 51.71% 62.25%
$R_{\text{ToolRL}}$ 75.27% 64.25%
$R_{\text{ToolRM}}$ 78.40% 64.32%
Qwen2.5- 3B-Instruct Base 43.06% 55.66%
$R_{\text{Schema}}$ 63.17% 66.54%
$R_{\text{ToolRL}}$ 80.42% 67.21%
$R_{\text{ToolRM}}$ 79.58% 67.51%
Table 5: BFCL-v3 performance of Llama-3.2-3B-Instruct and Qwen2.5-3B-Instruct trained with GRPO under three reward designs.
We train Llama-3.2-3B-Instruct and Qwen2.5-3B-Instruct models using the three reward variants and evaluate them on the BFCL-v3 dataset, with results shown in Table 5. Across both models, all three reward variants substantially improve performance over the base models. Simple schema-based rewards provide strong gains without requiring ground-truth supervision, yielding average improvements of about 20 points. ToolRM-based rewards achieve the best overall results: for Llama-3.2-3B-Instruct, $R_{\text{ToolRM}}$ attains the highest accuracy on both evaluation metrics, and for Qwen2.5-3B-Instruct, it provides the best Live AST accuracy, surpassing the gold-dependent $R_{\text{ToolRL}}$. Overall, $R_{\text{ToolRM}}$ consistently matches or exceeds $R_{\text{ToolRL}}$ despite not requiring access to ground truth, highlighting its practicality and effectiveness as a scalable reward signal for reinforcement learning in the tool-calling setting. Additional experimental details are provided in Appendix A.8.
@@ 6 Conclusion
In this paper, we presented a comprehensive framework for reward modeling in tool-calling scenarios. Our benchmark, FC-RewardBench, enables systematic assessment of reward models on tool-calling tasks. We also presented a framework for training outcome RMs that outperform existing significantly larger RMs in the tool calling setting. Overall, ToolRM enables effective inference-time scaling, data-efficient fine-tuning, and ground-truth–free policy optimization. Looking ahead, we see several promising directions for advancing reward modeling in this domain. First, moving beyond classification-based RMs to generative verifiers with chain-of-thought reasoning could improve robustness and interpretability. Second, incorporating the tool and environment state into training could help models safely recover from execution failures. Finally, bridging outcome and process reward modeling may offer a unified framework that balances scalability with fine-grained control over reasoning quality.
@@ Appendix A Appendix
Dataset # Examples # Tools (avg./query) # MT queries Avg. MT turns Nested calls Avg. output tool calls Data source
BFCL-v3 4,441 2,631 (3.3) 800 4.2 2.4 Real
API-Bank 473 64 (3.4) 397 3.4 x 1.0 Real
ToolAlpaca 100 64 (5.6) 0 x 1.5 Synthetic
NexusRaven 318 65 (7.4) 0 x 1.0 Synthetic
SealTools 627 3,036 (9.9) 0 2.9 Synthetic
Table 6: Statistics of the evaluation benchmarks. “MT” denotes multi-turn queries.
@@ A.1 FC-RewardBench benchmark details
A representative example from the FC-RewardBench dataset is shown in Figure 3.
Figure 3: Representative example from FC-RewardBench  the parameter player_count is set to an incorrect value. The tool catalog is hidden for brevity.
@@ A.4.1 Model Details
Additionally, we include the Tool-Augmented Reward Model (ernie-research/Themis-7b).
@@ A.4.2 LLM-as-Judge Prompt
A.5 Correlation with performance on downstream tasks:
Figure 5: Correlation heatmap between FC-RewardBench performance and downstream accuracy across generator models and benchmarks, showing consistently strong alignment (avg. correlation = 0.84).
A.6 Error Analysis
@@
A.7 ToolRM generalization
To assess how well ToolRM generalizes to inputs unseen during training, we embed each tool and conversation from both the training set and the five non-BFCL test sets using the all-mpnet-base-v2555https://huggingface.co/sentence-transformers/all-mpnet-base-v2 embedding model. We then compute, for each test example, the maximum cosine similarity with the training set and plot the resulting distributions in Figure 6. We observe low similarity between training and test examples for both tools and conversations (median << 0.6 in both cases), indicating that ToolRM generalizes effectively to novel inputs at test time.
Figure 6: Histogram of maximum cosine similarity between train and test tools and conversations.
A.8 GRPO Experiment Details
We utilize the open-source library verl [^6] https://github.com/volcengine/verl and the code base and dataset provided by ToolRL  [^7] https://github.com/qiancheng0/ToolRL to train models for GRPO experiments. The hyperparameters used to train the model are listed in Table 9, and the prompt used to train the models is listed in Listing 3.
Hyperparameter Value
Max prompt length 2048
Max response length 1024
Learning rate 1e-6
PPO mini batch size 128
Number of rollouts 4
Total epochs 15
Train batch size 512
Validation batch size 128
Table 9: GRPO Experiment Hyperparameters
1You are a helpful multi-turn dialogue assistant capable of leveraging tool calls to solve user tasks and provide structured chat responses.
2
3**Available Tools**
4“‘json
5{{TOOLS}}
6“‘
7
8**Steps for Each Turn**
91. **Think:** Recall relevant context and analyze the current user goal.
102. **Decide on Tool Usage:** If a tool is needed, specify the tool and its parameters.
113. **Respond Appropriately:** If a response is needed, generate one while maintaining consistency across user queries.
12
13**Output Format**
14“‘plaintext
15<think> Your thoughts and reasoning </think>
16<tool_call>
17{"name": "Tool name", "parameters": {"Parameter name": "Parameter content", "… …": "… …"}}
18{"name": "… …", "parameters": {"… …": "… …", "… …": "… …"}}
19
20</tool_call>
21<response> AIs final response </response>
22“‘
23
24**Important Notes**
251. You must always include the ‘<think>‘ field to outline your reasoning. Provide at least one of ‘<tool_call>‘ or ‘<response>‘. Decide whether to use ‘<tool_call>‘ (possibly multiple times), ‘<response>‘, or both.
262. You can invoke multiple tool calls simultaneously in the ‘<tool_call>‘ fields. Each tool call should be a JSON object with a "name" field and an "parameters" field containing a dictionary of parameters. If no parameters are needed, leave the "parameters" field an empty dictionary.
273. Refer to the previous dialogue records in the history, including the users queries, previous ‘<tool_call>‘, ‘<response>‘, and any tool feedback noted as ‘<obs>‘ (if exists).
Listing 3: Prompt used for GRPO training of models
A.9 Ablation analysis
We conduct an ablation study to assess the contribution of individual components to the overall performance of ToolRM. Specifically, we train ToolRM-1.5B, ablating different hyperparameters, training datasets, and generator models, and evaluating each variant on the FC-RewardBench dataset. Table 10 summarizes the results. First, we observe that the full ToolRM-1.5B model, incorporating all components, achieves the highest performance. Second, removing the API-Gen dataset leads to a 16.4-point drop in performance, underscoring its significance in training. Finally, obfuscating tool and parameter names results in a 13.63-point reduction in performance. This suggests that obfuscation prevents the model from overfitting to specific tool or parameter names and encourages it to attend to other parts of the tool specifications, thereby improving robustness and generalization.
Model Accuracy
ToolRM-1.5B 81.88%
Hyperparameter Ablation
Without obfuscation 68.25%
High reward centering ($\eta=0.1$) 80.02%
No reward centering ($\eta=0$) 80.32%
Data Ablation
Without API-Gen 65.48%
Without SGD 81.03%
Generator model ablation
Only large (>=>=12B) models 78.70%
Only small (<=<=2B) models 81.01%
Table 10: Ablation results for ToolRM-1.5B model on FC-RewardBench dataset.